• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name bogdanw

    Log entry time 00:13:37 on November 5,2006

    Entry number 188897

    keyword=shift summary

    continue data taking for B10 kinematics
    reboot IOC NMR for HRS-L and HRS-R
    find about 10% reduced number of hits in FPP
    it was this way but stable since Monday
    alarm in gas shed - one CO2 bank is empty
    take runs 1956-1965